OpenStudy (anonymous):

Trig help, please. Medal and fan if you help me out. 1. Evaluate cos(a+b) sin a = -7/25, cot b = -8/15 neither are in quadrant IV 2. Evaluate cos(a-b) tan a = 4/3, sin b = -7/25; pi < a < 3pi/2, 3pi/2 < b < 2pi (Exact answers no decimals)

OpenStudy (anonymous):

Here are some hints 1.)If you are giving 1 trigonometric ratio, you can find all the others using identities 2.)Try using the formula cos(a+b)=cosacosb-sinasinb 3.)The pythagorean triplets in questions are 7,24,25 and 8,15,17 and 3,4,5

OpenStudy (anonymous):

\(\huge\color{red}{cos(a+b)=cosacosb-sinasinb}\) \(\huge\color{red}{cos(a-b)=cosacosb+sinasinb}\)

OpenStudy (anonymous):

Thank you guys so much this really helped

OpenStudy (anonymous):

:)